# BCS-BEC crossover of the strongly interacting 6Li–40K mixture

### v1

@@ III BCS-BEC crossover
The fit to our QMC results gives ${\xi=0.3726(6)}$, ${\zeta=0.900(2)}$ and ${\nu=0.46(1)}$. The values of $\zeta$ and $\nu$ are very similar to the case of equal masses reported in Ref. Gandolfi et al. (2011). We also find a good qualitative agreement with the calculation of Ref. Diener and Randeria (2010).
The various tests we performed indicate that around ${1/a\,k_{F}\approx 0}$ the extrapolation to $r_{e}\rightarrow 0$ just shifts the energy but does not change the curvature of the equation of state in the region where $11/|a\,k_{F}|<$ 1. There, only the $\xi$ parameter is particularly sensitive to $r_{e}$ as is also the case for the equal mass system Gandolfi et al. (2011). It is interesting to compare the finite range extrapolation for the unequal mass mixture to the equal mass case. The comparison is motivated for the 6Li–40K system as the resonance is more narrow than for equal masses Naik et al. (2011) and consequently the unitary limit could be ”less universal” and the effect of the effective range more important. We performed several calculations at unitarity for different values of $r_{e}$ in order to test the finite range effects and extrapolated to zero effective range as shown in Fig. 3. The QMC points have been fit using the linear function
@@
and the best estimate given by QMC is $\xi=0.368(1)$. The value of $c$ gives the slope of the extrapolation that is $c=0.13(1)$ for unequal masses. It is rather intriguing that the slope of the energy as a function of the effective range is very similar to the equal mass case Carlson et al. (2011).
Figure 3: (color online) The energy of the 6Li–40K mixture as a function of the effective range $r_{e}$ of the interaction in the unitary limit. The energy is in units of the Fermi gas energy $E_{FG}$. The fit parameters are very similar to those found in the equal mass case in Gandolfi et al. (2011). The details of the fit to the QMC results are discussed in the main text.
@@
Figure 4: (color online) Heavy-light pair distribution function for different two-body interactions indicated in the legend. From left to right they correspond to $41/ak_{F}=-$ 4, $-1$, $-0.5$, $-0.1$, 00, $0.1$ and $0.5$. In the inset we show the same function multiplied by $(k_{F}r)^{2}$ for selected interaction strengths corresponding to $41/ak_{F}=-$ 4, $-1$, $-0.5$, 00 and $0.5$ (from the bottom to the top).
Figure 5: (color online) Momentum distribution $n(k)$ for different interaction strengths (same color convention as Fig. 4). In the inset we plot $(k/k_{F})^{4}n(k)$ to show the tail behavior of the momentum distribution. The dashed black line shows the contact parameter $C$ derived from the equation of state around the unitary limit Werner and Castin (2012).
@@
We have also computed the pair distribution function $g_{h-l}(r)$ between heavy and light particles, as shown in Fig. 4. $g_{h-l}(r)$ is calculated by a mixed estimate of the form Gezerlis and Carlson (2010),
$g_{h-l}(r)=\sum_{i<j^{\prime}}\left\langle\Psi_{0}\right|\delta(r_{ij}-r)O_{ij}^{P}\left|\Psi_{T}\right\rangle,$ (10)
where $\Psi_{0}$ is the DMC projected ground state.
@@
The function $g_{h-l}(r)$ exhibits a large peak at short distances, induced by the strong attractive interaction in the s-wave channel. In the weak coupling limit the value of the peak is smaller as expected, and is clear in the figure. In the inset of Fig. 4 we show $(k_{F}r)^{2}g_{h-l}(r)$ that can be used to extract the contact parameter as shown in Ref. Gandolfi et al. (2011). The limit of $(k_{F}r)^{2}g_{h-l}(r)$ for short distances approaches zero in the BCS limit as the short–range correlations are weaker when the strength of the attractive interaction decreases. On the other side, in the strong coupling regime where $0.51/ak_{F}=$ 0.5, the limit of $(k_{F}r)^{2}g_{h-l}(r)$ at small distances rapidly increases.
@@
In Fig. 5 we show the momentum distribution of the mass imbalanced system, calculated as the Fourier transform of the one-body density matrix by,
@@
where the integral over the solid angle is handled stochastically and the integral over $\delta r=|\bm{r}^{\prime}_{n}-\bm{r}_{n}|$ is carried out on a line to avoid statistical errors Gezerlis and Carlson (2010). In the very weak coupling region with $41/ak_{F}=-$ 4 the $n(k)$ is very similar to the non interacting Fermi gas momentum distribution, where $n(k)=1$ for $k<k_{F}$ and zero otherwise. Moving to strong couplings the tail of $n(k)$ is longer, as expected, due to the strong pairing effects. This is well evident in the inset of Fig. 5 where the function $(k/k_{F})^{4}n(k)$ is shown. The inset clearly shows the $C/k^{4}$ universal behavior predicted by Tan Tan_2008a; Tan (2008a, b), where $C$ is the contact parameter. For $01/ak_{F}=$ 0 the black dashed line shows the value of the contact derived from the equation of state Werner and Castin (2012) for Fig. 1 and Eq. (8).
IV Quasiparticle spectrum
@@
The quasiparticle dispersion is computed by adding an unpaired particle to the system Carlson et al. (2003); Chang et al. (2004); Gezerlis et al. (2009); Gandolfi et al. (2008); Gezerlis and Carlson (2010). We have carried out simulations of 67 particles where the unpaired particle is characterized by a finite momentum $|k|$ and mass $m_{l}$ or $m_{h}$. We then combine the result with the energy of the fully paired unpolarized system of 66 particles, and define the quasiparticle energy as
@@
where $k$ is the momentum of the extra particle, and $E(67)$ and $E(66)$ are the total energies of the systems with 67 and 66 particles. The two simulations have been performed at the same total density. In Ref. Gezerlis et al. (2009) the quasiparticle dispersion was computed by comparing the energy of the two systems at the same background density, then at constant volume. Here we have fixed the total density in order to avoid small effects due to the effective range of the interaction. However, the finite size effects should be very small as the volume of the simulation box is quite large.
The results of the quasiparticle dispersion in Fig. 6 show that the dispersion of light and heavy particles are very different from what has already been found at unitarity Gezerlis et al. (2009). For each case the quasiparticle dispersion is lower on the BCS side and increases by crossing unitarity and moving to the BEC side. In particular for light quasiparticle the position of the minimum changes from about $(k/k_{F})^{2}=0.5$ at $ak_{F}=-2$ to $(k/k_{F})^{2}=0$ at $ak_{F}=2$.
V Conclusions
In this paper we show QMC results exploring the BCS-BEC crossover of the mass imbalanced 6Li–40K Fermi mixture. QMC techniques have been used in the past to give the best upperbound of the energy of strongly interacting fermions in the unitary limit. We made use of QMC to compute the equation of state as a function of the two–body scattering length, studying the crossover from the weakly interacting limit where the system is well described by the BCS theory, to the strong coupling limit where fermions form molecules, and the system is well described by a weakly repulsive Bose gas. We have carefully checked finite-range effects close to the unitary limit, and provided an accurate fit of the equation of state based on our QMC calculations. We also show other properties of the unpolarized superfluid system, namely the pair distribution function and the momentum distribution. Finally, by considering the system with very small polarizations, we have computed the heavy- and light-quasiparticle dispersion for different strengths of the interaction. These results should provide a comprehensive framework for future theoretical and experimental investigations into the mass-imbalanced Fermi gas across the full range of the BCS-BCS crossover.
@@ Acknowledgments
The work of S.G. is supported by the U.S. Department of Energy, Office of Nuclear Physics, under contract No. DE-AC52-06NA25396, and by the Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) NUCLEI program. The work of R.C. and A.G. was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada and the Canada Foundation for Innovation (CFI). R.C. was also supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20220541ECR. Computational resources have been provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001. Computer time was also made available by the National Energy Research Scientic Computing Center (NERSC), and SHARCNET through the Digital Research Alliance of Canada.

### v2

@@ III BCS-BEC crossover
Figure 3: (color online) The energy of the 6Li-40K mixture as a function of the effective range $r_{e}$ of the interaction in the unitary limit. The energy is in units of the Fermi gas energy $E_{FG}$. The slope of the extrapolation is similar to what was found in the equal-mass case with lattice calculations Carlson et al. (2011). The details of the fit to the QMC results are discussed in the text.
The fit to our QMC results near unitarity gives ${\xi=0.3726(6)}$, ${\zeta=0.900(2)}$, and ${\nu=0.46(1)}$. Our result for $\xi$ differs appreciably from the equal-mass case near unitarity ($\xi=0.383(1)$ from Gandolfi et al. (2011)), which is expected to be a result of non-zero total momentum pairs in the ground state Gezerlis et al. (2009). However, the values of $\zeta$ and $\nu$ are very similar to what was found for the equal-mass case Ref. Gandolfi et al. (2011). Our results are in good qualitative agreement with the calculation of Ref. Diener and Randeria (2010), deviating from mean-field calculations which predict no difference between the equal- and unequal-mass cases.
The various tests we performed indicate that around ${1/a\,k_{F}\approx 0}$ the extrapolation to $r_{e}\rightarrow 0$ just shifts the energy and does not change the curvature of the equation of state in the region where $11/|a\,k_{F}|<$ 1. There, only the $\xi$ parameter is particularly sensitive to $r_{e}$, which is also the case for the equal-mass system Gandolfi et al. (2011). It is interesting to compare the finite-range extrapolation for the unequal-mass mixture to the equal-mass case. The comparison is warranted for the 6Li-40K system because the resonance is narrower than for equal masses Naik et al. (2011), and consequently, the unitary limit could be “less universal”, and the effect of the effective range could be more important. We performed several calculations at unitarity for different values of $r_{e}$ in order to test the finite-range effects and extrapolated to zero effective range as shown in Fig. 3. The QMC results were fit using the linear function
@@
and the best estimate given by QMC is $\xi=0.368(1)$. Similar to the above case near unitarity, our value for $\xi$ at the unitarity limit is slightly lower than what is found in equal-mass calculations carried out on the lattice ($\xi=0.372(3)$ from Carlson et al. (2011). This is further evidence of the contribution of nonzero total momentum pairs in the ground state, as found in Gezerlis et al. (2009). The value of $c=0.13(1)$ gives the slope of the extrapolation, and it is rather intriguing that the slope of the energy as a function of the effective range is very similar to the equal-mass case Carlson et al. (2011). This fact, along with the similarities to higher order terms from Eq. (9), should be explored in future investigations into mass-imbalanced cold atomic systems.
As discussed above, there are many reasons to expect different physics between the unequal-mass Fermi gas that we study and the more common equal-mass case. At the level of the equation of state the two systems display certain similarities (higher-order terms in the equation of state), as well as differences (the Bertsch parameter $\xi$ is quantitatively different near and at the unitary limit). To investigate further similarities and differences between the equal- and unequal-mass Fermi gases, we consider a number of different physical observables.
@@
Figure 4: (color online) Heavy-light pair-distribution function for different two-body interactions indicated in the legend. From left to right they correspond to $41/ak_{F}=-$ 4, $-1$, $-0.5$, $-0.1$, 00, $0.1$, and $0.5$. In the inset we show the same function multiplied by $(k_{F}r)^{2}$ for selected interaction strengths corresponding to $41/ak_{F}=-$ 4, $-1$, $-0.5$, 00, and $0.5$ (from the bottom to the top).
@@
We first consider the pair-distribution function $g(r)$ between heavy and light particles, as shown in Fig. 4. $g(r)$ is typically calculated by a mixed estimate of the form
$g(r)=\sum_{i<j^{\prime}}\frac{\left\langle\Psi_{0}\right|\delta(r_{ij^{\prime}}-r)O_{ij^{\prime}}^{P}\left|\Psi_{T}\right\rangle}{\left\langle\Psi_{0}|\Psi_{T}\right\rangle},$ (11)
where $\Psi_{0}$ is the DMC projected ground state and, for our purposes ($s$-wave central interaction) the $O_{ij^{\prime}}^{p}$ operator is equal to 1 Gezerlis and Carlson (2010). The sum is only over pairs of particles with different masses.
@@
The function $g(r)$ exhibits a large peak at short distances, induced by the strong attractive interaction in the $s$-wave channel. In the weak-coupling limit the value of the peak is smaller as expected, and is clear in our results. In the inset of Fig. 4 we show $(k_{F}r)^{2}g(r)$, which can be used to extract the contact parameter, as shown in Ref. Gandolfi et al. (2011). The limit of $(k_{F}r)^{2}g(r)$ for short distances approaches zero in the BCS limit as the short-range correlations are weaker when the strength of the attractive interaction decreases. On the other hand, in the strong-coupling regime where $0.51/ak_{F}=$ 0.5, the limit of $(k_{F}r)^{2}g(r)$ at small distances rapidly increases. Our results for unitarity (red diamonds in Fig. 4) can be compared with the equal-mass case shown in Fig. 4 of Gandolfi et al. (2011). One notable difference is that in the equal-mass case $g(r)$ decays with increasing $k_{F}r$, and we find this effect to be dramatically reduced for the unequal-mass case. By looking at the inset of Fig. 4, at unitarity we see that $(k_{F}r)^{2}g(r)$ is essentially constant over the considered range of $k_{F}r$. This can be interpreted as the heavy-light pairs remaining farther apart from one another, potentially due to the higher inertia of the heavy particles, compared to the equal-mass case.
Figure 5: (color online) Momentum distribution $n(k)$ for the unequal-mass system at unitarity. We calculate $n(k)$ using only the light particles (solid line) and using only the heavy particles (dashed line). In the inset we plot $(k/k_{F})^{4}n(k)$ to show the tail behavior of the momentum distribution. The dotted black line shows the contact parameter $C$ derived from our calculated equation of state around the unitary limit, following the prescription from Werner and Castin (2012)
Figure 6: (color online) Momentum distribution $n(k)$ for different interaction strengths (same color convention as in Fig. 4), using only the light particles. In the inset we plot $(k/k_{F})^{4}n(k)$ to show the tail behavior of the momentum distribution. The results for unitarity (red diamonds) also appear in Fig. 5.
@@
In Figs. 5 and 6 we show momentum distributions for the mass imbalanced system, calculated as the Fourier transform of the one-body density matrix by,
@@
where the integral over the solid angle is handled stochastically and the integral over $\delta r=|\bm{r}^{\prime}_{n}-\bm{r}_{n}|$ is carried out on a line to avoid statistical errors Gezerlis and Carlson (2010). The primed coordinate refers to a single particle that is moved along this line in order to carry out the integration over $d\delta r$. Equation (12) can then be computed for each particle in our system, and an average value can be taken. In Fig. 5 we first investigate the momentum distribution for the system at unitarity, averaging over only the light or heavy particles independently. We also investigate the tail behavior of the momentum distribution by plotting $(k/k_{F})^{4}n(k)$. We find in both cases a value for Tan’s contact parameter that agrees with the expected result from Werner and Castin (2012), given our equation of state in Fig. 1. Figure 5 also highlights a subtle difference between the momentum distributions of the heavy and light particles, with the heavy-particle calculations showing a noticeable decrease for low-momentum values. This feature is absent in studies of the equal-mass Fermi gas.
@@
Since considering either the light or heavy particles alone gives the same value for the contact, we can investigate the crossover behavior of the momentum distribution, as shown in Fig. 6, considering only the light particles. In the very weak coupling region with $41/ak_{F}=-$ 4, $n(k)$ is very similar to the non interacting Fermi gas momentum distribution, where $n(k)=1$ for $k<k_{F}$ and zero otherwise. Moving to strong couplings the tail of $n(k)$ is longer, as expected, due to the strong pairing effects. This is clearly evident in the inset of Fig. 6 where the function $(k/k_{F})^{4}n(k)$ is shown. The inset clearly shows the $C/k^{4}$ universal behavior predicted by Tan Tan (2008a, b, c), where $C$ is the contact, regardless of interaction strength. The value of $C$ we find from the tail of $n(k)$ agrees with the value that can be computed from a derivative of the equation of state with respect to the scattering length Werner and Castin (2012). It is also interesting to note that the value of $C$ we find for the unequal-mass case agrees within error with what is found for the equal-mass case Gandolfi et al. (2011). This highlights a fundamental similarity between the equal- and unequal-mass Fermi gases.
Last, the quasiparticle dispersion is computed by adding an unpaired particle to the system Carlson et al. (2003); Chang et al. (2004); Gezerlis et al. (2009); Gandolfi et al. (2008); Gezerlis and Carlson (2010). We carried out simulations of 67 particles where the unpaired particle is characterized by a finite momentum $|k|$ and mass $m_{l}$ or $m_{h}$. We then combine the result with the energy of the fully paired unpolarized system of 66 particles and define the quasiparticle energy as
@@
where $k$ is the momentum of the extra particle, and $E(67)$ and $E(66)$ are the total energies of the systems with 67 and 66 particles. The two simulations were performed at the same total density. In Ref. Gezerlis et al. (2009) the quasiparticle dispersion was computed by comparing the energy of the two systems at the same background density, then at constant volume. Here we have fixed the total density in order to avoid small effects due to the effective range of the interaction. However, the finite-size effects should be very small as the volume of the simulation box is quite large.
The results of the quasiparticle dispersion in Fig. 7 show that the dispersions of light and heavy particles are very different from what was already found at unitarity Gezerlis et al. (2009). For each case the quasiparticle dispersion is lower on the BCS side and increases by crossing unitarity and moving to the BEC side. In particular for light quasiparticles the position of the minimum changes from about $(k/k_{F})^{2}=0.5$ at $ak_{F}=-2$ to $(k/k_{F})^{2}=0$ at $ak_{F}=2$. Our results at unitarity agree with previous findings Gezerlis et al. (2009) that the quasiparticle dispersion for the unequal-mass Fermi gas differs dramatically from the equal-mass case Carlson and Reddy (2005). We have also shown that the differences in the quasiparticle dispersions for heavy and light particles are seen throughout the BCS-BEC crossover region.
IV Conclusions
In this paper we showed QMC results exploring the BCS-BEC crossover of the mass-imbalanced 6Li-40K Fermi mixture. QMC techniques were used in the past to give the best upper bound of the energy of strongly interacting fermions in the unitary limit. We made use of QMC to compute the equation of state as a function of the two-body scattering length, studying the crossover from the weakly interacting limit where the system is well described by the BCS theory to the strong coupling limit where fermions form molecules, and the system is well described by a weakly repulsive Bose gas. We carefully checked finite-range effects close to the unitary limit and provided an accurate fit of the equation of state based on our QMC calculations. We also showed other properties of the unpolarized superfluid system, namely, the pair-distribution function and the momentum distribution. Finally, by considering a system with very small polarizations, we computed the heavy- and light-quasiparticle dispersions for different strengths of the interaction. Throughout we also paid careful attention to the differences and similarities between the equal- and unequal-mass Fermi gases, finding several important results. Our calculations of the equation of state revealed that the higher order terms in the equation of state are very similar. We also found that the universal contact parameter is identical to that in the equal-mass case. However, we did find several key differences from the equal-mass case, such as the Bertsch parameter, quasiparticle dispersion, and tail of the pair-distribution function. Our results should provide a comprehensive framework for future theoretical and experimental investigations into mass-imbalanced Fermi gases across the full range of the BCS-BCS crossover.
@@ Acknowledgments
The work of S.G. is supported by the U.S. Department of Energy, Office of Nuclear Physics, under Contract No. DE-AC52-06NA25396, and by the Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) NUCLEI program. The work of R.C. and A.G. was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada and the Canada Foundation for Innovation (CFI). R.C. was also supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under Project No. 20220541ECR. Computational resources were provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001. Computer time was also made available by the National Energy Research Scientific Computing Center (NERSC) and SHARCNET through the Digital Research Alliance of Canada.